A line has a slope of 6 and passes through the point (12, 5). What is the equation of the line?

Answer by ewatrrr(24785)


Hi
line with a slope of 6 and passing through the point (12, 5).


y = 6x + b |using ordered pair Pt(12,5) to solve for b
5 = 6*12 + b
5 = 72 + b
-67 = b
y = 6x - 67 |equation in standard slope-intercept form